Perform the indicated operation and, if possible, simplify. Please show all of your work so I know how to do it.
3/14+2/5

Answer by stanbon(75887)

3/14+2/5
----
lcd = 70
---
Rewrite each fraction with the lcd as its denominator:
-----
= (3*5)/lcd + (2*14)/lcd
----
Combine the numerators over the lcd:
= (15+28)/lcd
-----
= 43/70
===================
Cheers,
Stan H.